OpenStudy (anonymous):

Find the x and y intercepts and then sketch the graph for: y=x^2 + 6x + 9

OpenStudy (anonymous):

what do u do for finding the intercepts on x and y axis?

OpenStudy (anonymous):

suppose u have to sketch 2x+3y=6

OpenStudy (anonymous):

find out, where the curve y=x^2 + 6x + 9 will cut the axes?

OpenStudy (anonymous):

yes i did this too find the x intercept but it doesn't make sense: y=(x+3)(x+3) <------ factorised let y = 0 x+3= 0 x=-3 but then both x intercepts = -3. that doesnt make sense

OpenStudy (anonymous):

for x intercept...put y=0 then x=-3 for y intercept ...put x=0 then y=9

OpenStudy (anonymous):

but then both x intercepts = -3. that doesnt make sense???

OpenStudy (anonymous):

there should be two different x intercept but they both =-3.

OpenStudy (anonymous):

u have to sketch y=(x+3)^2 It'll cut the axes at (-3,0) and (0,9)

OpenStudy (anonymous):

its a parabola

OpenStudy (anonymous):

for x intercept, we put y=0 so (x+3)(x+3)=0

OpenStudy (anonymous):

then we get a repeated root x=-3...and it'll cut the x-axis once only

hartnn (hartnn):

it makes sense that this graph only has 1 x-intercept, because it touches x-axis only at a point (-3,0) , that is x axis is tangent to this graph.....